# Gradient Aligned Regression via Pairwise Losses

### v3

@@ B.8 Experiment Settings for AgeDB
We follow the the same experiment settings from the previous work (Zha et al., 2023). The age range is between 0 and 101. It is split into 12208/2140/2140 images for training/validation/testing sets. The SGD optimizer and cosine learning rate annealing is utilized for training (Loshchilov and Hutter, 2016). The batch size is set to 256. For both training from scratch and linear probe experiments, we select the best learning rates and weight decays for each dataset by grid search, with a grid of learning rates from {0.01, 0.05, 0.1, 0.2, 0.5, 1.0} and weight decays from {1e-6, 1e-5, 1e-4, 1e-3}. For the predictor training of two-stage methods, we adopt the same search setting as above except for adding additional 0 weight decay to the search choices of weight decays. The RNC temperature parameter is search from {0.1, 0.2, 0.5, 1.0, 2.0, 5.0} and select the best, which is 2.0. We train all one-stage methods and the encoder of two-stage methods for 400 epochs, and the linear regressor of two-stage methods for 100 epochs. The hyper-parameters for all the compared methods are searched in the same region as described in the main content. For training from scratch experiments, because the ConR (Keramati et al., 2023) is proposed to use two data augmentations, we utilize two data augmentations for each training data for all compared methods. For each method, we randomly repeat five times with different random seeds to report the mean and standard deviation for evaluations.

### v4

@@ B.8 Experiment Settings for AgeDB
We follow the the same experiment settings from the previous work (Zha et al., 2023). The age range is between 0 and 101. It is split into 12208/2140/2140 images for training/validation/testing sets. The SGD optimizer and cosine learning rate annealing is utilized for training (Loshchilov and Hutter, 2016). The batch size is set to 256. For both training from scratch and linear probe experiments, we select the best learning rates and weight decays for each dataset by grid search, with a grid of learning rates from {0.01, 0.05, 0.1, 0.2, 0.5, 1.0} and weight decays from {1e-6, 1e-5, 1e-4, 1e-3}. For the predictor training of two-stage methods, we adopt the same search setting as above except for adding additional 0 weight decay to the search choices of weight decays. The RNC temperature parameter is search from {0.1, 0.2, 0.5, 1.0, 2.0, 5.0} and select the best, which is 2.0. We train all one-stage methods and the encoder of two-stage methods for 400 epochs, and the linear regressor of two-stage methods for 100 epochs. The hyper-parameters for all the compared methods are searched in the same region as described in appendix B.7. For training from scratch experiments, because the ConR (Keramati et al., 2023) is proposed to use two data augmentations, we utilize two data augmentations for each training data for all compared methods. The $\alpha$ for GAR is selected as 0.1 for training from scratch setting and as 1.0 for linear probe setting. For each method, we randomly repeat five times with different random seeds to report the mean and standard deviation for evaluations.
@@ B.10 Sensitivity Analysis
B.11 Different Batch Sizes
We conduct experiments on different batch sizes {32, 64, 128, 256, 512} for GAR. The experimental setting is kept as the same as the main experiments in Appendix B.7. The results are included in Tab. 6. Because the proposed GAR learns both the mean and variance of the label values by conventional loss and proposed pairwise losses (by Theorem 1), we additionally provide the convergences curves for the estimated means and standard deviations in Fig. 12. From both the table and the figures, we observe a strong correlation between more accurate estimation of mean and standard deviation to better testing performance (Batch Size = 32 on Concrete Compressive Strength and 64 on Parkinson Motor).
Table 6: Testing performance with different batch sizes.
Dataset Concrete Compressive Strength Parkinson(Total) Parkinson(Motor)
Batch Size MAE RMSE Pearson Spearman MAE RMSE Pearson Spearman MAE RMSE Pearson Spearman
32 3.866(0.132) 5.625(0.194) 0.942(0.005) 0.944(0.004) 1.875(0.306) 3.957(1.684) 0.925(0.063) 0.944(0.014) 1.518(0.074) 2.661(0.201) 0.944(0.008) 0.942(0.007)
64 3.966(0.183) 5.705(0.257) 0.939(0.005) 0.937(0.009) 1.431(0.109) 2.737(0.182) 0.967(0.004) 0.961(0.007) 1.184(0.033) 2.321(0.098) 0.958(0.004) 0.954(0.004)
128 4.167(0.113) 5.837(0.178) 0.937(0.004) 0.938(0.002) 1.61(0.075) 2.971(0.12) 0.961(0.003) 0.954(0.003) 1.315(0.06) 2.455(0.097) 0.953(0.004) 0.949(0.005)
256 4.603(0.075) 6.222(0.148) 0.929(0.004) 0.931(0.003) 2.64(0.071) 4.258(0.206) 0.922(0.009) 0.907(0.01) 1.652(0.069) 2.68(0.134) 0.943(0.006) 0.939(0.006)
512 4.989(0.139) 6.568(0.226) 0.917(0.006) 0.921(0.008) 4.032(0.056) 5.679(0.04) 0.855(0.007) 0.83(0.007) 2.709(0.062) 3.849(0.126) 0.891(0.016) 0.883(0.018)
Figure 12: Batch Sizes on mean and standard deviation estimation for GAR
B.12 Difference with Standardization
The formulation for $\mathcal{L}_{\text{diff}}^{\text{MSE}}$ and $\mathcal{L}_{\text{diffnorm}}^{p=2}$ in Eq. 5 and Eq. 3 might remind reader of data centering or standardization. In this subsection, we highlight the difference:
  • $\mathcal{L}_{\text{diff}}^{\text{MSE}}$ and $\mathcal{L}_{\text{diffnorm}}^{p=2}$ ‘centralize’ or ‘standardize’ not only for ground truth label, but also for prediction. Besides, the predictions are dynamic during model training process.
  • The proposed GAR, Alg. 1, learns the mean and standard deviation dynamically for each mini-batch during training, rather than for the whole dataset.
  • We tested all the baselines on the five tabular datasets with standardization on both data features and labels, where the results are summarized in Tab. 7. From the results, the proposed GAR enjoys similar superiority on the standardized datasets as the previous main results in Tab. 1.
Table 7: Testing performance on the standardized tabular datasets/settings. Mean and standard deviation are reported. ‘Gains’ includes the percentages that GAR outperform the MAE and the best competitor (in parenthesis). The ‘p-values’ are calculated based on Student’s t-test for the ‘Gains’. p-value < 0.05 usually means significant difference. $\downarrow$ means the smaller the better; $\uparrow$ means the larger the better.
Standardized Dataset Method MAE MSE Huber Focal (MAE) Focal (MSE) RankSim RNC ConR GAR (ours) Gains (%) p-values
Concrete Compressive Strength MAE $\downarrow$ 0.284(0.003) 0.288(0.002) 0.303(0.004) 0.289(0.005) 0.299(0.01) 0.802(0.012) 0.283(0.007) 0.286(0.007) 0.278(0.003) 2.22(1.77) 0.02(0.24)
RMSE $\downarrow$ 0.39(0.02) 0.387(0.009) 0.396(0.01) 0.386(0.008) 0.391(0.014) 0.994(0.005) 0.38(0.002) 0.378(0.006) 0.38(0.011) 2.56(-0.66) 0.41(0.7)
Pearson $\uparrow$ 0.924(0.003) 0.919(0.005) 0.918(0.002) 0.921(0.004) 0.921(0.004) 0.674(0.059) 0.925(0.003) 0.925(0.002) 0.925(0.004) 0.11(0.0) 0.68(0.93)
Spearman $\uparrow$ 0.927(0.003) 0.927(0.002) 0.919(0.004) 0.922(0.005) 0.924(0.003) 0.706(0.029) 0.929(0.002) 0.927(0.005) 0.929(0.002) 0.28(0.0) 0.16(0.97)
Wine Quality MAE $\downarrow$ 0.574(0.003) 0.611(0.006) 0.615(0.003) 0.59(0.001) 0.62(0.005) 0.595(0.005) 0.616(0.005) 0.579(0.005) 0.565(0.007) 1.56(1.56) 0.05(0.05)
RMSE $\downarrow$ 0.824(0.005) 0.787(0.009) 0.79(0.005) 0.794(0.005) 0.79(0.007) 0.817(0.009) 0.81(0.003) 0.811(0.006) 0.79(0.004) 4.17(0.0) 0.0(0.53)
Pearson $\uparrow$ 0.566(0.004) 0.614(0.011) 0.612(0.006) 0.605(0.005) 0.611(0.009) 0.571(0.011) 0.587(0.002) 0.586(0.007) 0.618(0.003) 9.22(0.67) 0.0(0.5)
Spearman $\uparrow$ 0.603(0.007) 0.63(0.009) 0.63(0.006) 0.624(0.003) 0.623(0.006) 0.603(0.006) 0.609(0.005) 0.616(0.005) 0.638(0.007) 5.85(1.18) 0.0(0.23)
Parkinson (Total) MAE $\downarrow$ 0.33(0.016) 0.321(0.01) 0.363(0.002) 0.323(0.013) 0.333(0.009) 0.807(0.001) 0.306(0.006) 0.354(0.004) 0.241(0.009) 26.81(21.12) 0.0(0.0)
RMSE $\downarrow$ 0.507(0.013) 0.473(0.012) 0.515(0.003) 0.505(0.011) 0.487(0.014) 10.828(19.644) 0.474(0.015) 0.524(0.006) 0.385(0.02) 24.14(18.73) 0.0(0.0)
Pearson $\uparrow$ 0.864(0.007) 0.883(0.006) 0.859(0.002) 0.866(0.008) 0.875(0.008) 0.203(0.052) 0.881(0.008) 0.853(0.003) 0.924(0.008) 6.91(4.63) 0.0(0.0)
Spearman $\uparrow$ 0.841(0.007) 0.86(0.006) 0.835(0.003) 0.845(0.006) 0.852(0.006) 0.188(0.043) 0.861(0.009) 0.826(0.009) 0.909(0.008) 8.05(5.57) 0.0(0.0)
Parkinson (Motor) MAE $\downarrow$ 0.309(0.017) 0.313(0.017) 0.362(0.004) 0.314(0.01) 0.339(0.02) 0.847(0.002) 0.313(0.01) 0.339(0.011) 0.225(0.014) 27.13(27.13) 0.0(0.0)
RMSE $\downarrow$ 0.487(0.028) 0.46(0.032) 0.514(0.004) 0.495(0.01) 0.486(0.026) 0.993(0.003) 0.477(0.014) 0.512(0.011) 0.37(0.02) 24.05(19.58) 0.0(0.0)
Pearson $\uparrow$ 0.871(0.015) 0.886(0.016) 0.856(0.002) 0.868(0.005) 0.872(0.014) 0.13(0.022) 0.878(0.007) 0.857(0.007) 0.928(0.008) 6.51(4.72) 0.0(0.0)
Spearman $\uparrow$ 0.864(0.016) 0.872(0.018) 0.847(0.001) 0.861(0.006) 0.865(0.015) 0.133(0.046) 0.866(0.007) 0.847(0.006) 0.923(0.009) 6.79(5.81) 0.0(0.0)
Super Conductivity MAE $\downarrow$ 0.244(0.001) 0.256(0.0) 0.263(0.001) 0.247(0.001) 0.264(0.002) 0.574(0.068) 0.241(0.001) 0.243(0.001) 0.223(0.002) 8.85(7.59) 0.0(0.0)
RMSE $\downarrow$ 0.396(0.002) 0.389(0.001) 0.399(0.001) 0.394(0.003) 0.392(0.002) 0.762(0.083) 0.385(0.001) 0.394(0.001) 0.358(0.003) 9.68(7.02) 0.0(0.0)
Pearson $\uparrow$ 0.919(0.0) 0.921(0.001) 0.917(0.0) 0.92(0.001) 0.92(0.001) 0.811(0.008) 0.923(0.001) 0.92(0.0) 0.934(0.001) 1.66(1.14) 0.0(0.0)
Spearman $\uparrow$ 0.916(0.001) 0.907(0.001) 0.901(0.001) 0.912(0.001) 0.9(0.002) 0.825(0.004) 0.916(0.001) 0.917(0.001) 0.928(0.001) 1.25(1.17) 0.0(0.0)
IC50 MAE $\downarrow$ 0.82(0.002) 0.829(0.003) 0.821(0.002) 0.821(0.002) 0.831(0.002) 0.82(0.002) 0.827(0.002) 0.814(0.001) 0.79(0.017) 3.7(2.97) 0.01(0.02)
RMSE $\downarrow$ 1.026(0.002) 1.016(0.001) 1.026(0.002) 1.025(0.002) 1.021(0.002) 1.031(0.005) 1.026(0.002) 1.024(0.003) 0.992(0.024) 3.33(2.35) 0.02(0.08)
Pearson $\uparrow$ 0.079(0.025) 0.166(0.027) 0.122(0.031) 0.089(0.019) 0.156(0.021) 0.015(0.025) -0.012(0.017) 0.224(0.031) 0.3(0.035) 278.81(33.82) 0.0(0.01)
Spearman $\uparrow$ 0.13(0.031) 0.165(0.032) 0.12(0.028) 0.11(0.02) 0.146(0.021) 0.026(0.041) 0.019(0.042) 0.198(0.024) 0.277(0.027) 112.94(39.9) 0.0(0.0)